OpenStudy (anonymous):

Please help me, this problem is driving me crazy. Simplify 4 over the quantity of negative 2 minus 2i.

OpenStudy (anonymous):

\[4 \over -2-2i\]

OpenStudy (anonymous):

@rainstorm12

OpenStudy (welshfella):

multiply top and bottom of the fraction by the complement of -2 - 2i which is -2 + 2i this removes the imaginary part of the complex number from the denominator

OpenStudy (anonymous):

ok, what is 2i multiplied by 2i btw?

OpenStudy (welshfella):

i^2 = -1

OpenStudy (anonymous):

4- 2i^2 = -8 + 8i

OpenStudy (anonymous):

is that right so far?

OpenStudy (welshfella):

thats correct for the numerator

OpenStudy (welshfella):

(-2 - 21)(-2 + 2i) = ?

OpenStudy (welshfella):

that should be an i not a 1

OpenStudy (anonymous):

idk what 2i multplied by 2i is

OpenStudy (welshfella):

its -2i * 2i -2 * 2 = -4 and i^2 = -1

OpenStudy (welshfella):

so it equals -1 *-4

OpenStudy (welshfella):

by definition i = sqrt(-1) so i^2 = -1

OpenStudy (anonymous):

i keep coming to a solution of i = 1

OpenStudy (welshfella):